Exercise 29.4
Undervalued assets, full and partial goodwill method
On 1 July 2022, Jane Ltd acquired 75% of the issued shares of Austen Ltd for $360 000. At this date, the equity of Austen Ltd consisted of share capital of $200 000 and retained earnings of $130 000. All the identifiable assets and liabilities of Austen Ltd were recorded at amounts equal to fair value except for the following.

Austen Ltd also had an internally generated patent not recognised at 1 July 2022. Jane Ltd assessed the fair value of that patent at $50 000. The tax rate is 30%.

Required

  1. Prepare the acquisition analysis at 1 July 2022 assuming that Jane Ltd used the partial

    goodwill method.

  2. Prepare the acquisition analysis at 1 July 2022 assuming that Jane Ltd used the full

    goodwill method and the fair value of the non‐controlling interest at 1 July 2022 was

    $110 000.

(LO4 and LO7)
Answer #1

1)

Book value of net assets of Austen ltd = 200000 + 130000 = 330000

Fair value = 330000 + 50000( asset not recognized) = 380000

Since the acquirer Jane ltd holds 75% of the shares, it owns , 75%.x 380000 = 285000

Goodwill= cash paid - acquirers share = 360000 - 285000 = 75000

Thus, Jane ltd will record the following

Net assets 380000

Goodwill 75000

To Cash 360000

To Non Controlling interest 95000

2)

Fair value of net assets = 330000 + 50000( asset not recognized) = 380000

Non-controlling interest is calculated as (goodwill under full goodwill method + assets acquired − liabilities assumed − cash paid)

110000 = goodwill + 380000-360000

goodwill = 90000

Thus, Jane ltd will record the following

Net assets 380000

Goodwill 90000

To Cash 360000

To Non Controlling interest 110000